OpenStudy (anonymous):

The function f(x) = 245(4)^x represents the growth of a fruit fly population every year in the orange grove. Stacey wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Stacey's purposes?

OpenStudy (anonymous):

f(x) = 45(4)^x f(x) = 45(4^12)^x/12 f(x) = 2752(4)x f(x) = 245(4 ^1/2)^12x

OpenStudy (anonymous):

@dan815